Decide whether the graph of the equation is a line, a parabola, or a circle.
x^2+4x+y^2-6y=3

Answer by stanbon(48581)

x^2+4x+y^2-6y=3
Complete the square on the x-terms and the y-terms separately to get:
(x^2+4x+4) + (y^2-6y+9) = 3+4+9
(x-2)^2 + (y-3)^2 = 16
Circle with center at (2,3) and radius of 4
=============================================
Cheers,
Stan H.